# Prompt-based Pseudo-labeling Strategy for Sample-Efficient Semi-Supervised Extractive Summarization

### v2

@@ Abstract
Semi-supervised learning (SSL) is a widely used technique in scenarios where labeled data is scarce and unlabeled data is abundant. While SSL is popular for image and text classification, it is relatively underexplored for the task of extractive text summarization. Standard SSL methods follow a teacher-student paradigm to first train a classification model and then use the classifier’s confidence values to select pseudo-labels for the subsequent training cycle; however, such classifiers are not suitable to measure the accuracy of pseudo-labels as they lack specific tuning for evaluation, which leads to confidence values that fail to capture the semantics and correctness of the generated summary. To address this problem, we propose a prompt-based pseudo-labeling strategy with LLMs that picks unlabeled examples with more accurate pseudo-labels than using just the classifier’s probability outputs. Our approach also includes a relabeling mechanism that improves the quality of pseudo-labels. We evaluate our method on three text summarization datasets: TweetSumm, WikiHow, and ArXiv/PubMed. We empirically show that a prompting-based LLM that scores and generates pseudo-labels outperforms existing SSL methods on ROUGE-1, ROUGE-2, and ROUGE-L scores on all the datasets. Furthermore, our method achieves competitive G-Eval scores (evaluation with GPT-4) as a fully supervised method that uses 100% of the labeled data with only 16.67% of the labeled data.
@@
Gaurav Sahu University of Waterloo gaurav.sahu@uwaterloo.ca Olga Vechtomova University of Waterloo Issam H. Laradji ServiceNow Research University of British Columbia
@@ 1 Introduction
Text summarization is a crucial task in today’s data-driven era. Major enterprises store customer service chatlogs and wish to create an automated text summarization system to enhance the user experience Feigenblat et al. (2021). With the exponential growth of digital content, the need for automated summarization techniques that swiftly process and distill information has become paramount Koupaee and Wang (2018); Cohan et al. (2018). Text summarization systems are broadly categorized into two types: abstractive, where the summaries are concise paraphrases of the original text (Barzilay and McKeown, 2005; Nallapati et al., 2016), and extractive, where the summaries combine existing sentences verbatim from the original text (Radev et al., 2004; Li et al., 2006; Wong et al., 2008). While both systems have their advantages, in today’s fast-paced lifestyle, extractive summarization systems have become increasingly relevant as they are fast and interpretable while preserving the source integrity and reducing potential misinterpretation.
@@
Standard deep learning-based methods reformulate extractive summarization as a binary text classification problem: given a sentence in the input document, predict its likelihood of appearing in the output summary (Amini and Gallinari, 2001; Wong et al., 2008). SSL techniques, in particular, follow a teacher-student paradigm to train a classifier on the limited labeled data first to generate pseudo-labels for the unlabeled set and then use the same classifier’s confidence values to select the best pseudo-labels for the next training cycle. Most recently, Liu and Lapata (2019) and Zhuang et al. (2023) employed this strategy using large language models (LLMs) like BERT (Devlin et al., 2019) as their backbones. While these works have shown promising results, the underlying classification model is not suitable to determine the accuracy of pseudo-labels because it lacks specific fine-tuning for measuring pseudo-label accuracy. On the other hand, recent instruction-tuned LLMs like GPT-4 (OpenAI, 2023) and LLaMA-2 (Touvron et al., 2023) have been successfully used for generating pseudo-labels and evaluation for a wide range of tasks Sahu et al. (2022, 2023); Dziri et al. (2023); Mishra et al. (2023); Liu et al. (2023); Zhang et al. (2023a).
Motivated by the success, this work uses a generative LLM like GPT-4 to improve the process of pseudo-labeling during the training phase of an SSL algorithm. Scoring all the pseudo-labels with LLMs can quickly get expensive as the number of pseudo-labels increases, so we first use the classification model’s confidence values to shortlist a fixed number of pseudo-labels. Then, we prompt GPT-4 with specific instructions to score the pseudo-labels and select top- pseudo-labels to include in the subsequent training cycle. Our method also includes a prompt-based relabeling mechanism that improves pseudo-labels’ accuracy.
We evaluate our approach on three text summarization datasets: TweetSumm (Feigenblat et al., 2021), WikiHow (Koupaee and Wang, 2018), and ArXiv/PubMed Cohan et al. (2018). We show that our method enhances the performance of a standard teacher-student pipeline to achieve competitive performance to that of a fully supervised model while using 6 less labeled data. Figure 1 further demonstrates the effectiveness of our method. In summary, we 1) propose a prompt-based pseudo-label selection strategy for semi-supervised extractive text summarization. 2) We also propose a relabeling mechanism using an LLM (such as GPT-4) to generate high-quality summaries for unlabeled examples. 3) Finally, we show that our method is more sample-efficient and outperforms prior semi-supervised methods for extractive summarization on ROUGE-1, ROUGE-2, ROUGE-L, and G-Eval scores across three benchmark datasets.
@@ 2 Related Work
The earliest works on extractive text summarization use unsupervised and supervised algorithms to extract sentence and word-level features. Some early examples of unsupervised approaches include using graph-based ranking techniques like LexRank (Erkan and Radev, 2004) and external knowledge from Wikipedia (Sankarasubramaniam et al., 2014), and fuzzy-logic systems (Kyoomarsi et al., 2008; Suanmali et al., 2009). While these methods improve coherency and reduce redundancy, they rely on surface-level features and cannot handle dangling anaphoras. Our method combines richer sentence representations of pretrained language models like BERT and DistilBERT Sanh et al. (2019) with GPT-4, which has recently been shown to help in anaphora resolution (Shen and Elhoseiny, 2023).
Prior supervised approaches use probabilistic support vector machines (Wong et al., 2008), Bayes rule (Neto et al., 2002), conditional random fields (Svore et al., 2007), and more recently, pretrained language models like BERT (Liu and Lapata, 2019; Pilault et al., 2020; Zhong et al., 2020) for the task. While supervised methods provide better representation for sentence selection, they need large human-annotated datasets for training, which can be financially expensive to gather. To address this problem, Zhuang et al. (2023) explore the semi-supervised setting and propose a novel objective function to improve the pretraining of BERT for dialog summarization. Their method significantly improves over the traditional pretraining process; however, it lacks flexibility and is heavily tailored for dialog summarization. Our method uses a generic prompt-based strategy with GPT-4 that can be easily adapted across different domains. Additionally, our method uses less data and generates higher-quality pseudo-labels.
@@ 3.2 Stage 2: Generating Pseudo-labels using the Teacher Model
We use the teacher model to generate pseudo-labels for the unlabeled set . Next, we shortlist a subset of 50 pseudo-labels with the highest teacher confidence. We compute the teacher confidence for a generated summary (a.k.a pseudo-label) as follows: let denote the probability with which the -th sentence in an unlabeled document is present in its summary , and let denote the indicator function defined as defined in Equation 1.
We then compute the teacher confidence for the pseudo-label by averaging the probabilities of selected sentences. Formally, we define the teacher confidence () for an input text as follows:
@@ 3.3 Stage 3: LLM Relabeling of Teacher’s Pseudo-labels
After selecting the top 50 pseudo-labels using teacher confidence defined in Equation 2, we prompt an LLM (GPT-4, in our case) to generate a summary for each shortlisted unlabeled example. This effectively relabels the pseudo-label from Stage 2. Specifically, we follow the prompt template in Figure 3 when generating summaries using GPT-4, which uses the same mechanism as PreSumm, i.e., we instruct GPT-4 to output probabilities for each sentence in the input document and then concatenate the top- lines in their order of appearance in the input text.
@@ 3.4 Stage 4: LLM Scoring of Pseudo-labels
In the last stage of our pipeline, we prompt GPT-4 as shown in Figure 4 to output a rating between 0-100 for the pseudo-labels from Stage 3. Specifically, we list the following criteria for evaluating the pseudo-labels:
@@
Here, is a hyperparameter chosen based on the average summary length in the dataset. Finally, we choose the top 5 pseudo-labeled examples with the highest GPT-4 scores and append them to the existing labeled set for the next training cycle.
We repeat Stages 1-4 times to improve the initial summarization model and use the model obtained after the last cycle for generating summaries for the unseen test set. Algorithm 1 also provides an overview of the proposed semi-supervised approach. To note, denotes generating summaries for using , and denotes the prompt-based relabeling of pseudo-labels from Stage 2. Furthermore, denotes scoring pseudo-labels (based on teacher confidence in Stage 2 and GPT-4 in Stage 4) and selecting the top pseudo-labels.
@@
| BERT | |||||||||
@@
4 Experimental Setup
4.1 Datasets
We use three datasets for extractive summarization.
1) TweetSumm Feigenblat et al. (2021) is a real-world customer service chat dataset. The dataset has 1100 conversations between a customer and an agent, and each conversation has three human-annotated extractive summaries. The training set has 858 dialogs, out of which we randomly sample a subset of 50 examples as the initial training set and treat the remaining as the unlabeled set. We use the official validation and test sets for evaluation.
Table 1 summarizes the dataset statistics. Notably, the WikiHow and ArXiv/PubMed datasets do not have extractive labels. Therefore, we follow the same steps as the original PreSumm paper (Liu and Lapata, 2019) to obtain the ground truth extractive labels. Additionally, due to the large training corpus size for WikiHow and ArXiv/PubMed, we limit the pool of unlabeled examples by randomly sampling a subset of 1,000 articles in each training cycle.
We choose the above three datasets to cover a diverse set of scenarios: TweetSumm contains real-world conversations and human-annotated extractive summaries, WikiHow contains medium-sized documents, and ArXiv/PubMed contains long documents and summaries.

### v3

@@ Abstract
Semi-supervised learning (SSL) is a widely used technique in scenarios where labeled data is scarce and unlabeled data is abundant. While SSL is popular for image and text classification, it is relatively underexplored for the task of extractive text summarization. Standard SSL methods follow a teacher-student paradigm to first train a classification model and then use the classifier’s confidence values to select pseudo-labels for the subsequent training cycle; however, such classifiers are not suitable to measure the accuracy of pseudo-labels as they lack specific tuning for evaluation, which leads to confidence values that fail to capture the semantics and correctness of the generated summary. To address this problem, we propose a prompt-based pseudo-labeling strategy with LLMs that picks unlabeled examples with more accurate pseudo-labels than using just the classifier’s probability outputs. Our approach also includes a relabeling mechanism that improves the quality of pseudo-labels. We evaluate our method on three text summarization datasets: TweetSumm, WikiHow, and ArXiv/PubMed. We empirically show that a prompting-based LLM that scores and generates pseudo-labels outperforms existing SSL methods on ROUGE-1, ROUGE-2, and ROUGE-L scores on all the datasets. Furthermore, our method achieves competitive L-Eval scores (evaluation with LLaMa-3) as a fully supervised method in a data-scarce setting and outperforms fully supervised method in a data-abundant setting.
@@
Gaurav Sahu & Olga Vechtomova David R. Cheriton School of Computer Science University of Waterloo, Canada {gaurav.sahu}@uwaterloo.ca Issam H. Laradji ServiceNow Research University of British Columbia
@@ 1 Introduction
Text summarization is a crucial task in today’s data-driven era. Major enterprises store customer service chatlogs and wish to create an automated text summarization system to enhance the user experience (Feigenblat et al., 2021). With the exponential growth of digital content, the need for automated summarization techniques that swiftly process and distill information has become paramount (Koupaee and Wang, 2018; Cohan et al., 2018). Text summarization systems are broadly categorized into two types: abstractive, where the summaries are concise paraphrases of the original text (Barzilay and McKeown, 2005; Nallapati et al., 2016), and extractive, where the summaries combine existing sentences verbatim from the original text (Radev et al., 2004; Li et al., 2006; Wong et al., 2008). While both systems have their advantages, in today’s fast-paced lifestyle, extractive summarization systems have become increasingly relevant as they are fast and interpretable while preserving the source integrity and reducing potential misinterpretation.
@@
Standard deep learning-based methods reformulate extractive summarization as a binary text classification problem: given a sentence in the input document, predict its likelihood of appearing in the output summary (Amini and Gallinari, 2001; Wong et al., 2008). SSL techniques, in particular, follow a teacher-student paradigm to train a classifier on the limited labeled data first to generate pseudo-labels for the unlabeled set and then use the same classifier’s confidence values to select the best pseudo-labels for the next training cycle. Most recently, Liu and Lapata (2019) and Zhuang et al. (2023) employed this strategy using large language models (LLMs) like BERT (Devlin et al., 2019) as their backbones. While these works have shown promising results, the underlying classification model is not suitable to determine the accuracy of pseudo-labels because it lacks specific fine-tuning for measuring pseudo-label accuracy. On the other hand, recent instruction-tuned LLMs like GPT-4 (OpenAI, 2023) and LLaMA-2 and LLaMa-3 (Touvron et al., 2023) have been successfully used for generating pseudo-labels and evaluation for a wide range of tasks (Sahu et al., 2022, 2023; Dziri et al., 2023; Mishra et al., 2023; Liu et al., 2023; Zhang et al., 2023a).
Motivated by the success, this work uses a generative LLMs to improve the process of pseudo-labeling during the training phase of an SSL algorithm. Scoring all the pseudo-labels with LLMs can quickly get expensive as the number of pseudo-labels increases, so we first use the classification model’s confidence values to shortlist a fixed number of pseudo-labels. Then, we prompt an LLM with specific instructions to score the pseudo-labels and select top- pseudo-labels to include in the subsequent training cycle. Our method also includes a prompt-based relabeling mechanism that improves pseudo-labels’ accuracy. We evaluate our approach on three text summarization datasets: TweetSumm (Feigenblat et al., 2021), WikiHow (Koupaee and Wang, 2018), and ArXiv/PubMed (Cohan et al., 2018). We show that our method enhances the performance of a standard teacher-student pipeline to achieve competitive performance to that of a fully supervised model while using 6 less labeled data in a data-scarce setting and outperforms fully supervised models in data-abundant setting. Figure 1 further demonstrates the effectiveness of our method.
In summary, we 1) propose a prompt-based pseudo-label selection strategy for semi-supervised extractive text summarization. 2) We also propose a relabeling mechanism using an LLM (such as GPT-4) to generate high-quality summaries for unlabeled examples. 3) Finally, we show that our method is more sample-efficient and outperforms prior semi-supervised methods for extractive summarization on ROUGE-1, ROUGE-2, ROUGE-L, and L-Eval (same as L-Eval but using LLaMA-3 as the LLM) scores across three benchmark datasets.
@@ 2 Related Work
The earliest works on extractive text summarization use unsupervised and supervised algorithms to extract sentence and word-level features. Examples of unsupervised approaches include using graph-based ranking techniques like LexRank (Erkan and Radev, 2004) and external knowledge from Wikipedia (Sankarasubramaniam et al., 2014), and fuzzy-logic systems (Kyoomarsi et al., 2008; Suanmali et al., 2009). While these methods improve coherency and reduce redundancy, they rely on surface-level features and cannot handle dangling anaphoras. Our method combines richer sentence representations of pretrained language models like BERT and DistilBERT (Sanh et al., 2019) with GPT-4, which has recently been shown to help in anaphora resolution (Shen and Elhoseiny, 2023).
Prior supervised approaches use probabilistic support vector machines (Wong et al., 2008), Bayes rule (Neto et al., 2002), conditional random fields (Svore et al., 2007), and more recently, pretrained language models like BERT (Liu and Lapata, 2019; Pilault et al., 2020; Zhong et al., 2020) for the task. While supervised methods provide better representation for sentence selection, they need large human-annotated datasets for training, which can be financially expensive to gather. To address this problem, Zhuang et al. (2023) explore the semi-supervised setting and propose a novel objective function to improve the pretraining of BERT for dialog summarization. Their method significantly improves over the traditional pretraining process; however, it lacks flexibility and is heavily tailored for dialog summarization. Our method uses a generic prompt-based strategy with LLMs that can be easily adapted across different domains. Additionally, our method uses less data and generates higher-quality pseudo-labels.
@@ 3.2 Stage 2: Generating Pseudo-labels using the Teacher Model
We use the teacher model to generate pseudo-labels for the unlabeled set . Next, we shortlist a subset of 50 pseudo-labels with the highest teacher confidence 111We experiment 5, 10, 20, 50, and 75 and find 50 to be optimal.. We compute the teacher confidence for a generated summary (a.k.a pseudo-label) as follows: let denote the probability with which the -th sentence in an unlabeled document is present in its summary , and let denote the indicator function: We then compute the teacher confidence for the pseudo-label by averaging the probabilities of selected sentences. We define the teacher confidence () for an input text as follows:
@@ 3.3 Stage 3: LLM Relabeling of Teacher’s Pseudo-labels
After selecting the top 50 pseudo-labels using teacher confidence defined in Equation 1, we prompt an LLM (GPT-4/LLaMA, in our case) to generate a summary for each shortlisted unlabeled example. This effectively relabels the pseudo-label from Stage 2. Specifically, we follow the prompt template in Figure 3(a) when generating summaries using GPT-4, which uses the same mechanism as PreSumm, i.e., we instruct the LLM to output probabilities for each sentence in the input document and then concatenate the top- lines in their order of appearance in the input text.
@@ 3.4 Stage 4: LLM Scoring of Pseudo-labels
In the last stage of our pipeline, we prompt LLaMA-3 as shown in Figure 3(b) to output a rating between 0-100 for the pseudo-labels from Stage 3. Specifically, we list the following criteria for evaluating the pseudo-labels:
@@
Here, is a hyperparameter chosen based on the average summary length in the dataset. Finally, we choose the top 5 pseudo-labeled examples with the highest LLaMA scores and append them to the existing labeled set for the next training cycle.
We repeat Stages 1-4 times to improve the initial summarization model and use the model obtained after the last cycle for generating summaries for the unseen test set. Algorithm 1 also provides an overview of the proposed semi-supervised approach. To note, denotes generating summaries for using , and denotes the prompt-based relabeling of pseudo-labels from Stage 2. Furthermore, denotes scoring pseudo-labels (based on teacher confidence in Stage 2 and LLM score in Stage 4) and selecting the top pseudo-labels.
4 Experimental Setup
4.1 Datasets
We use three datasets for extractive summarization.
1) TweetSumm (Feigenblat et al., 2021) is a real-world customer service chat dataset. The dataset has 1100 conversations between a customer and an agent, and each conversation has three human-annotated extractive summaries. The training set has 858 dialogs, out of which we randomly sample a subset of 50 examples as the initial training set and treat the remaining as the unlabeled set. We use the official validation and test sets for evaluation.
@@ 4.1 Datasets
Table 1 summarizes the dataset statistics. Notably, the WikiHow and ArXiv/PubMed datasets do not have extractive labels. Therefore, we follow the same steps as the original PreSumm paper (Liu and Lapata, 2019) to obtain the ground truth extractive labels. Additionally, due to the large training corpus size for WikiHow and ArXiv/PubMed, we limit the pool of unlabeled examples by randomly sampling a subset of 1,000 articles in each training cycle. We choose the above three datasets to cover a diverse set of scenarios: TweetSumm contains real-world conversations and human-annotated extractive summaries, WikiHow contains medium-sized documents, and ArXiv/PubMed contains long documents.
4.2 Training and Evaluation Details.
Training. We use the TransformerSum repository222https://transformersum.readthedocs.io/en/latest/ to implement our training pipeline. We use PreSumm as our teacher model and experiment with two backbones: distilbert-base-uncased and bert-base-uncased. We perform experiments in two settings: 1) data-scarce setting where fix the size of the labeled set to 50 for all the datasets, and 2) data-abundant setting where we set the size of to 500. We set to 50 for all experiments. We add 5 pseudo-labels to the training set in each cycle, thus resulting in a final training set size of 300 (50 labeled + 250 pseudo-labeled examples.) We set the summary size to 4 for TweetSumm and 8 for WikiHow and ArXiv/PubMed. We base these summary sizes on the average summary size of the labeled training set. For training, we start with a learning rate of on all the datasets and use a cyclic learning rate scheduler during training (Smith, 2015), which is the default setting in TransformerSum. Additionally, we use AdamW as our optimizer with . We train all our models on a single V100 GPU with 12G VRAM. We repeat each experiment for three different seeds and report the mean and standard deviation in our results unless otherwise stated.
@@
Confidence + L-3 score | 45.7 (1.1) | 36.9 (0.2) | 47.8 (1.2) | 21.6 (0.4) | 4.1 (0.5) | 11.1 (0.8) | 23.9 (0.9) | 6.1 (0.3) | 12.9 (1.3) |
+ L-3 relabeling (Ours) | 56.2 (1.1) | 45.1 (1.2) | 55.9 (1.1) | 22.3 (0.1) | 5.8 (0.2) | 13.6 (0.3) | 24.5 (0.6) | 7.7 (1.4) | 15.7 (0.3) |
BERTbase (50 labels) | |||||||||
@@
Confidence + L-3 score | 51.7 (1.2) | 41.6 (1.2) | 50.3 (1.2) | 25.9 (0.3) | 5.2 (0.2) | 13.0 (0.8) | 27.6 (0.4) | 7.9 (0.5) | 15.3 (1.1) |
+ L-3 relabeling (Ours) | 58.4 (1.2) | 50.1 (0.3) | 59.1 (1.2) | 26.0 (0.2) | 6.9 (0.3) | 15.1 (0.2) | 29.0 (0.5) | 9.4 (0.7) | 17.4 (0.3) |
BERTbase (500 labels) | |||||||||
TSL (500:500) | 59.0 | 48.3 | 58.2 | - | - | - | - | - | - |
Random | 55.1 (1.4) | 42.7 (1.1) | 50.3 (1.2) | 25.6 (1.3) | 4.5 (1.1) | 15.2 (1.3) | 25.4 (1.5) | 9.5 (1.2) | 24.1 (1.2) |
Confidence | 61.8 (0.7) | 54.9 (0.8) | 60.3 (0.9) | 28.4 (0.6) | 8.0 (1.1) | 22.5 (1.0) | 29.4 (0.5) | 11.5 (0.6) | 27.7 (0.8) |
+ L-3 score | 63.4 (0.5) | 55.6 (0.8) | 62.1 (0.7) | 28.9 (0.4) | 8.2 (0.5) | 28.4 (0.4) | 31.7 (0.3) | 11.8 (0.2) | 29.4 (0.4) |
+ L-3 relabeling (Ours) | 64.2 (0.2) | 56.2 (0.4) | 62.8 (0.6) | 30.7 (0.4) | 8.8 (0.3) | 29.5 (0.3) | 33.5 (0.3) | 12.3 (0.2) | 32.2 (0.3) |
Evaluation. First, we use the standard ROUGE- (R-) and ROUGE-L (R-L) F1 scores (Lin, 2004) on the test set to evaluate summary quality. ROUGE- measures the -gram overlap between the predicted and the ground truth summaries, whereas ROUGE-L also takes into consideration the order of -grams. We use pyrouge to compute ROUGE scores in our setup and report them in Table 2. Additionally, we use L-Eval, which is the same as G-Eval (Liu et al., 2023) but uses LLaMA instead of GPT-4 as the backend. We use a metric like G-Eval as it substantially surpasses ROUGE scores and previous model-based evaluation metrics, such as BERTScore (Zhang et al., 2019) and BARTScore (Yuan et al., 2021), for summarization. More importantly, it is shown to be better aligned with human preferences than the existing automated evaluation metrics for text summarization. Finally, G-Eval has been shown to not exhibit bias towards GPT-generated content. We report L-Eval results in Table 3.
4.3 Baselines
We compare our method with the following baselines: 1) PreSumm (Liu and Lapata, 2019). The original PreSumm model that pretrains a BERT model for summarization. We train two PreSumm models – one on a limited training set with 50 labeled examples to match the starting point of our semi-supervised setting and another with 300 labeled examples, the same as the dataset size at the end of our training cycle. 2) Teacher-Student Learning (TSL) (Zhuang et al., 2023). Current state-of-the-art semi-supervised method on TweetSumm. The teacher-student learning framework uses a similar formulation for computing model confidence to ours, as follows: . Here, , where is the probability of sentence being selected for summary for dialog estimated by the teacher model, and . We report the performance of the TSL (50:500) and TSL (500:500) models from the paper, as they are the closest to our setup (50/500 labeled examples + 500 unlabeled examples). 3) Confidence + G-4 relabeling + G-4 score (Ours). Our proposed method following the methodology in Section 3. We first use the PreSumm teacher model to shortlist 50 pseudo-labels (Stage 1 and 2), relabel them using GPT-4 (Stage 3), and then select the top 5 using GPT-4 score (Stage 4). 4) Confidence + G-4 score. We skip Stage 3 from 3) to directly score the top 50 PreSumm pseudo-labels using GPT-4. We run this baseline to measure the effect of relabeling in our pipeline. 5) Confidence + G-4 relabeling. We skip Stage 4 from 3) and select the final 5 pseudo-labels based on PreSumm confidence. 6) Confidence + L-3 relabeling + L-3 score (Ours). Same as 3) but using LLaMA-3. 7) Confidence + L-3 score (Ours). Same as 4) but using LLaMA-3. 8) Confidence. We skip Stage 3 and 4 from from 3) and select 5 PreSumm pseudo-labels based on PreSumm confidence. 9) Random. Same as 6) but instead of using teacher confidence defined in Equation 1, we randomly select five PreSumm pseudo-labels to include in each cycle.